# Clock Auctions Augmented with Unreliable Advice

### v1

@@
We provide the first analysis of clock auctions through the learning-augmented framework. Deferred-acceptance clock auctions are a compelling class of mechanisms satisfying a unique list of highly practical properties, including obvious strategy-proofness, transparency, and unconditional winner privacy, making them particularly well-suited for real-world applications. However, early work that evaluated their performance from a worst-case analysis standpoint concluded that no deterministic clock auction can achieve much better than an $O(\log(n))$ approximation of the optimal social welfare (where $n$ is the number of bidders participating in the auction), even in seemingly very simple settings.
To overcome this overly pessimistic impossibility result, which heavily depends on the assumption that the designer has no information regarding the preferences of the participating bidders, we leverage the learning-augmented framework. This framework assumes that the designer is provided with some advice regarding what the optimal solution may be. This advice may be the product of machine-learning algorithms applied to historical data, so it can provide very useful guidance, but it can also be highly unreliable.
Our main results are learning-augmented clock auctions that use this advice to achieve much stronger performance guarantees whenever the advice is accurate (known as consistency), while simultaneously maintaining worst-case guarantees even if this advice is arbitrarily inaccurate (known as robustness). Specifically, for the standard notion of consistency, we provide a clock auction that achieves the best of both worlds: $(1+\epsilon)$-consistency for any constant $\epsilon>0$ and $O(\log{n})$ robustness. We then also consider a much stronger notion of consistency and provide an auction that achieves the optimal trade-off between this notion of consistency and robustness.
@@ 1 Introduction
Our focus in this paper is on a classic family of mechanism design problems involving a set $N$ of $n$ buyers competing for some type of resource or service, and a seller (the service provider) who needs to decide which subset of the buyers to serve. Each buyer $i\in N$ has a value $v_{i}$ for receiving the service (i.e., the price they are willing to pay for it), and there is a feasibility constraint $\mathcal{F}\subseteq 2^{N}$ that restricts the sets of agents that can be simultaneously served. The goal of the service provider is to achieve an efficient outcome, i.e., to serve a feasible set of buyers $F\in\mathcal{F}$ that maximizes the social welfare: $\sum_{i\in F}v_{i}$. Depending on the nature of the feasibility constraint $\mathcal{F}$, this captures many well-studied optimization problems, like the knapsack problem, the maximum weight independent set problem, and many more. However, apart from the computational challenges of the underlying optimization problem, the seller also faces the additional obstacle that the values of the buyers are private (not known to the seller), making it even harder to maximize the social welfare.
@@
To address the fact that buyers’ values are private, a common solution in the mechanism design literature is to use direct revelation mechanisms: these mechanisms ask each buyer to report their value to the mechanism, which can then use this information to decide which buyers should receive the service and how much they should pay for it. However, unless these mechanisms are carefully designed, they are very likely to introduce incentives for sellers to misreport their true values (e.g., to report a lower value in order to reduce their payment for the service). A central goal in mechanism design is to design strategyproof mechanisms, which ensure that sellers’ optimal strategy is to always report their true values. A classic example of a strategyproof direct revelation mechanism that maximizes social welfare is the Vickrey-Clarke-Groves (VCG) mechanism. However, despite its elegance and great theoretical appeal, the VCG mechanism is rarely used in practice due to a variety of shortcomings [LonelyVCG]. For example, its strategyproofness heavily depends on its ability to solve the underlying optimization problem optimally, even though it may be NP-hard. In fact, direct revelation mechanisms in general face several issues that make them impractical for many applications. For example, (i) even if they are actually strategyproof, it may be non-trivial for buyers to verify it, (ii) they require that buyers put a lot of trust in the designer, and (iii) buyers need to directly reveal all of their private information, which can be sensitive and could potentially be used against them in the future (e.g., to set higher reserve prices).
@@
To overcome the issues that direct revelation mechanisms face, MS2014, MS2019 proposed deferred-acceptance clock auctions as a much more practical alternative for high-stakes auctions. In a clock auction, buyers are not asked to report their private values. Instead, the auction takes place over a sequence of rounds, and each buyer is offered a personalized price that weakly increases over time. Buyers can then remain active as long as they are willing to pay the price offered to them, and they can drop out whenever the price offered exceeds their value. These auctions have a wide variety of benefits that make them very practical: they guarantee (i) transparency (there is no way the auctioneer can mishandle the information provided by the buyers behind the scenes), (ii) unconditional winner privacy (the winners of the auction never need to reveal their true value), and (iii) simplicity (buyers do not need to understand the inner workings of the auction; all they need to know is the price offered to them in each round). These are properties that most strategyproof auctions do not satisfy.
@@
Motivated by this unique list of very appealing properties, subsequent work focused on analyzing the worst-case approximation guarantees that clock auctions can achieve for a variety of different feasibility constraints, setting aside computational constraints and focusing on information-theoretic constraints. The first set of results was rather pessimistic, showing that no deterministic clock auction can guarantee an approximation much better than $O(\log n)$ even in seemingly simple settings [DGR17]. On the positive side, subsequent work showed that there exists a natural clock auction, the water-filling clock auction (WFCA), that can guarantee an approximation of $O(\log n)$ for any type of feasibility constraint [CGS22]. Although the water-filling auction achieves the optimal worst-case approximation guarantee, a logarithmic approximation may not be quite as appealing from a practical standpoint. This is a common issue when analyzing algorithms and mechanisms from a worst-case perspective, leading to results that may be overly pessimistic.
@@
To overcome the overly pessimistic nature of worst-case analysis, recent work has seen a surge in using the learning-augmented framework, which assumes that the designer is provided with some useful, though unreliable, prediction or advice regarding the instance at hand. Using this advice, the goal is to provide stronger and much more practical performance guarantees whenever this advice is accurate (the consistency guarantee), while simultaneously maintaining non-trivial worst-case guarantees even if the advice is arbitrarily inaccurate (the robustness guarantee). As a result, this framework provides a natural way to leverage machine-learned predictions to guide the design of mechanisms while maintaining the important robustness that comes with worst-case analysis.
@@ 1.1 Our Results
In this paper, we provide the first analysis of clock auctions through the learning-augmented framework. Specifically, rather than assuming that the auction has absolutely no information regarding the values of the bidders, we assume that it is provided with a prediction $\hat{\texttt{OPT}}\in\mathcal{F}$ regarding which feasible set in $\mathcal{F}$ is the one that has the optimal social welfare. Crucially, this prediction can be arbitrarily inaccurate, so the auction needs to use it carefully to maintain any bounded robustness.
@@
In stark contrast to prior work, our first main result is a very optimistic one: we propose a learning-augmented clock auction that simultaneously achieves the “best of both worlds”. Specifically, our auction achieves a constant approximation of the optimal social welfare whenever the prediction is accurate, i.e., $\hat{\texttt{OPT}}$ is indeed optimal, while maintaining the worst-case approximation guarantee of $O(\log n)$, irrespective of how inaccurate the prediction may be. In fact, we show that it achieves a consistency of $1+\epsilon$ for an arbitrarily small constant $\epsilon>0$ while maintaining a robustness of $O(\frac{1}{\epsilon}\log n)$, and we prove that this is asymptotically optimal even with respect to its dependence on $\epsilon$. Finally, we also extend our auction to achieve a smoothly degrading approximation guarantee even when the prediction is not exactly accurate, up to some predetermined error tolerance.
@@
For our second main result we move beyond the standard notion of consistency, and consider a significantly stronger notion, which we denote as consistency. In contrast to the standard consistency constraint that binds only if the prediction is fully accurate, consistency binds on every input and requires that the social welfare achieved by the auction always approximates the social welfare of the predicted set, $\hat{\texttt{OPT}}$. Note that, whenever $\hat{\texttt{OPT}}$ is indeed optimal, approximating its social welfare reduces to approximating the optimal social welfare, so consistency is strictly more demanding than the standard notion of consistency. This notion of consistency dates back to the earlier work of MNS07 and has also been studied explicitly or implicitly in some subsequent work in the learning-augmented literature. Our main result for this notion is a learning-augmented clock auction that achieves an asymptotically optimal trade-off between robustness and consistency. Specifically, we show that the optimal robustness that can be achieved by a learning-augmented auction that satisfies $\alpha$-consistency, is $\Theta(n^{1/(\alpha-1)}\log{n})$.
@@ 1.2 Related Work
Clock auctions
@@
Since their formal introduction in [MS2014, MS2019], clock auctions have attracted attention from both the economics and computer science communities, partly due to the many practical advantages they offer over the sealed-bid format. These important practical properties include: (i) obvious strategyproofness, a strong notion of incentive compatibility originally defined by li2017obviously; (ii) unconditional winner privacy defined by MS2019 which guarantees that clock auctions reveal the minimum amount of information possible regarding the value of the winners; and (iii) transparency, including the fact that ascending clock auctions are considered "credible" auctions as defined in [akbarpour2020credible]. Notably, clock auctions are the unique class of auctions satisfying these properties, making them uniquely well-suited to practical implementation. We refer the reader to [MS2019] and [FGGS22, Section 1.2] for a more complete discussion of the properties of clock auctions.
@@
The papers technically closest to our work include [BLP2009, CGS22, DGR17]. Before [MS2014, MS2019] explicitly defined clock auctions, BLP2009 proposed a deterministic, prior-free clock auction that achieved a $O(\log{v_{\max}})$-approximation, where $v_{\max}$ is the largest value of any single bidder. However, their auction notably achieved only a $\Theta(n)$-approximation. In terms of lower bounds, DGR17 demonstrated that no deterministic, prior-free clock auction could achieve a $\log^{\tau}{n}$-approximation for any constant $\tau<1$. Over a decade after [BLP2009], CGS22 showed that the bound of [DGR17] was essentially tight by providing a deterministic, prior-free clock auction that achieved an $O(\log{n})$-approximation (and simultaneously an $O(\log{v_{\max}})$-approximation).
@@
In single-parameter, forward auction settings, CGS22 were the first to propose using randomization to circumvent the impossibility results of [DGR17]. They provided an auction guaranteeing expected welfare achieving an $O(\sqrt{\log{k}})$-approximation to the optimal social welfare, where $k$ is the number of maximal feasible sets in the set system $\mathcal{F}$. This was later improved to a $O(\log{\log{k}})$-approximation guarantee by [FGGS22]. Beyond the single-parameter, binary service level, forward auction context. For example, GMR17 demonstrated how to extend clock auctions to settings with multiple levels of service and analyzed several special cases of this general framework. Many works have also examined reverse (i.e., procurement) clock auctions, both with [JM17, BGGST22, ensthaler2014dynamic, BB2023, HHCT23] and without [K2015, bichler2020strategyproof] an auctioneer budget constraint. The works of [LM2020, DTR2017] studied double clock auctions involving both strategic sellers and buyers. GPPS21 studied clock auctions in the important economic setting of interdependent values. Finally, GH23 examined clock auctions for the purpose of consumer utility maximization.
Learning-augmented framework
@@
In recent years, the learning-augmented framework has gained widespread recognition as a valuable paradigm for designing and analyzing algorithms. We refer to [MV22] for a survey of early contributions and [alps] for a frequently updated list of papers in this field. This approach aims to overcome the limitations of overly pessimistic worst-case bounds. In just the past five years, over 200 papers have revisited classic algorithmic problems using this framework, with prominent examples including online paging [lykouris2018competitive], scheduling [PSK18], optimization problems involving covering [BMS20] and knapsack constraints [IKQP21], as well as Nash social welfare maximization [banerjee2020online], the secretary problem [AGKK23, DLLV21, KY23], and a range of graph-related challenges [azar2022online].As we point out in our results section, there is in fact another stronger notion of consistency that requires a good approximation of the predicted solution irrespective of the quality of predictions. This notion was first proposed in [MNS07], and subsequent work, including but not limited to [BMS20, KN23], also defines this as consistency. Additionally, we note that although some works do not explicitly state it, their results actually hold for the stronger notion of consistency [WZ20, JLLTZ22, ABGOT22].
@@
More closely related to our work, the line of research on learning-augmented mechanisms interacting with strategic agents is recent and was initiated by ABGOT22 and XL22. It encompasses strategic facility location [ABGOT22, XL22, IB22, BGT24], strategic scheduling [XL22, BGT223], auctions [MV17, XL22, LuWanZhang23, caragiannis2024randomized, BGTZ23], bicriteria mechanism design (optimizing both social welfare and revenue) [BPS23], graph problems with private input [CKST24], metric distortion [BFGT23], and equilibrium analysis [GKST22, IBB24]. Recently, CSV24 revisited mechanism design problems with predictions on the outcome space instead of the input. For further reading on this line of work, we refer to [BGT23].

### v2

@@
We provide the first analysis of (deferred acceptance) clock auctions in the learning-augmented framework. These auctions satisfy a unique list of very appealing properties, including obvious strategyproofness, transparency, and unconditional winner privacy, making them particularly well-suited for real-world applications. However, early work that evaluated their performance from a worst-case analysis perspective concluded that no deterministic clock auction with $n$ bidders can achieve a $O(\log^{1-\epsilon}n)$ approximation of the optimal social welfare for a constant $\epsilon>0$, even in very simple settings. This overly pessimistic impossibility result heavily depends on the assumption that the designer has no information regarding the bidders’ values. Leveraging the learning-augmented framework, we instead consider a designer equipped with some (machine-learned) advice regarding the optimal solution; this advice can provide useful guidance if accurate, but it may be unreliable.
Our main results are learning-augmented clock auctions that use this advice to achieve much stronger performance guarantees whenever the advice is accurate (known as consistency), while maintaining worst-case guarantees even if this advice is arbitrarily inaccurate (known as robustness). Our first clock auction achieves the best of both worlds: $(1+\epsilon)$-consistency for any desired constant $\epsilon>0$ and $O(\log{n})$ robustness; we also extend this auction to achieve error tolerance. We then consider a much stronger notion of consistency, which we refer to as consistency, and provide an auction that achieves a near-optimal trade-off between consistency and robustness. Finally, using our impossibility results regarding this trade-off, we prove lower bounds on the “cost of smoothness,” i.e., on the robustness that is achievable if we also require that the performance of the auction degrades smoothly as a function of the prediction error.
@@ 1 Introduction
Our focus in this paper is on a classic family of mechanism design problems involving a set $N$ of $n$ buyers competing for some type of resource or service, and a seller (the service provider) who needs to decide which subset of buyers to serve. Each buyer $i\in N$ has a value $v_{i}$ for receiving the service (i.e., the price they are willing to pay for it), and there is a feasibility constraint $\mathcal{F}\subseteq 2^{N}$ that restricts the sets of agents that can be served simultaneously. The goal of the service provider is to achieve an efficient outcome, i.e., to serve a feasible set of buyers $F\in\mathcal{F}$ that maximizes social welfare: $\sum_{i\in F}v_{i}$. Depending on the nature of the feasibility constraint $\mathcal{F}$, this captures many well-studied optimization problems, like the knapsack problem, the maximum weight independent set problem, and many more. However, apart from the computational challenges of the underlying optimization problem, the seller also faces the additional obstacle that the values of the buyers are private (not known to the seller), making it even harder to maximize social welfare.
@@
To address the fact that buyers’ values are private, a common solution in the mechanism design literature is to use direct revelation mechanisms: these mechanisms ask each buyer to report their value to the mechanism, which can then use this information to decide which buyers should receive the service and how much they should pay for it. However, unless these mechanisms are carefully designed, they are very likely to introduce incentives for buyers to misreport their true values (e.g., to report a lower value to reduce their payment for the service). A central goal in mechanism design is to design strategyproof mechanisms, which ensure that buyers’ optimal strategy is to always report their true values. A classic example of a strategyproof direct revelation mechanism that maximizes social welfare is the Vickrey-Clarke-Groves (VCG) mechanism. However, despite its elegance and great theoretical appeal, the VCG mechanism is rarely used in practice due to a variety of shortcomings [LonelyVCG]. For example, its strategyproofness heavily depends on its ability to solve the underlying optimization problem optimally, even though it may be NP-hard. In fact, direct revelation mechanisms in general face several issues that make them impractical for many applications. For example, (i) even if they are actually strategyproof, it may be non-trivial for buyers to verify it [KHL1987], (ii) they require that buyers put a lot of trust in the designer [akbarpour2020credible], and (iii) buyers need to directly reveal all of their private information, which can be sensitive and could potentially be used against them in the future (e.g., to set higher reserve prices) [rothkopf1990vickrey].
@@
To overcome the issues that direct revelation mechanisms face, Milgrom and Segal [MS2014, MS2019] proposed deferred-acceptance clock auctions as a much more practical alternative for high-stakes auctions. In a clock auction, buyers are not asked to report their private values. Instead, the auction takes place over a sequence of rounds, and each buyer is offered a personalized price that weakly increases over time. Buyers can then remain active as long as they are willing to pay the price offered to them, and they can drop out whenever the price offered exceeds their value. These auctions have a wide variety of benefits that make them very practical: they guarantee (i) transparency (there is no way the auctioneer can mishandle the information provided by the buyers behind the scenes), (ii) unconditional winner privacy (the winners of the auction never need to reveal their true value), and (iii) simplicity (buyers do not need to understand the inner workings of the auction; all they need to know is the price offered to them in each round). These are properties that most strategyproof auctions do not satisfy.
@@
Motivated by this unique list of very appealing properties, subsequent work focused on analyzing the worst-case approximation guarantees that clock auctions can achieve for a variety of different feasibility constraints, often setting aside computational constraints and focusing on information-theoretic constraints. The first set of results was rather pessimistic, showing that no deterministic clock auction can guarantee an approximation much better than $O(\log n)$ even in seemingly simple settings [DGR17]. On the positive side, subsequent work showed that there exists a natural clock auction, the water-filling clock auction (WFCA), that can guarantee an approximation of $O(\log n)$ for any type of feasibility constraint [CGS22]. Although the water-filling auction achieves the optimal worst-case approximation guarantee, a logarithmic approximation may not be quite as appealing from a practical standpoint. This is a common issue when analyzing algorithms and mechanisms from a worst-case perspective, leading to results that may be overly pessimistic.
@@
To overcome the overly pessimistic nature of worst-case analysis, a surge of recent work has used the learning-augmented framework, which assumes that the designer is provided with some useful, though unreliable, prediction or advice regarding the instance at hand. Using this advice, the goal is to provide stronger and much more practical performance guarantees whenever this advice is accurate (the consistency guarantee), while simultaneously maintaining non-trivial worst-case guarantees even if the advice is arbitrarily inaccurate (the robustness guarantee). As a result, this framework provides a natural way to leverage machine-learned predictions to guide the design of mechanisms while maintaining the important robustness that comes with worst-case analysis.
@@ 1.1 Our Results
We provide the first analysis of clock auctions within the learning-augmented framework. Specifically, rather than assuming that the auction has absolutely no information regarding the values of the bidders, we assume it is provided with a prediction, $\hat{\texttt{OPT}}\in\mathcal{F}$, regarding which feasible set in $\mathcal{F}$ has the optimal social welfare. Crucially, this prediction can be arbitrarily inaccurate, so the auction must use it carefully to maintain any bounded robustness.
@@
In stark contrast to prior work, our first main result is highly optimistic: we propose a learning-augmented clock auction that simultaneously achieves the "best of both worlds." Specifically, our auction achieves a constant approximation of the optimal social welfare whenever the prediction is accurate, i.e., $\hat{\texttt{OPT}}$ is indeed optimal, while maintaining the worst-case approximation guarantee of $O(\log n)$, regardless of how inaccurate the prediction may be. In fact, we show that it achieves a consistency of $1+\epsilon$ for any desired, arbitrarily small constant $\epsilon>0$, while maintaining a robustness of $O(\frac{1}{\epsilon}\log n)$, and we prove that this is asymptotically optimal even with respect to its dependence on $\epsilon$. We also extend this auction to achieve error tolerance, i.e., an approximation guarantee that degrades gracefully as a function of the error, up to some predetermined error tolerance threshold.
@@
For our second main result, we move beyond the standard notion of consistency and consider a significantly stronger notion, which we refer to as consistency. In contrast to the standard consistency constraint that binds only if the prediction is fully accurate, consistency binds on every input and requires that the social welfare achieved by the auction always approximates the social welfare of the predicted set, $\hat{\texttt{OPT}}$. Note that, whenever $\hat{\texttt{OPT}}$ is indeed optimal, approximating its social welfare reduces to approximating the optimal social welfare, so consistency is strictly more demanding than the standard notion of consistency.
This notion of consistency dates back to earlier work by Mahdian et al. [MNS07] and has also been studied explicitly or implicitly in some subsequent work in the learning-augmented literature. Our main result for this notion is a learning-augmented clock auction that achieves a near-optimal trade-off between robustness and consistency. Specifically, we show that the optimal robustness achievable by any learning-augmented clock auction that satisfies $\alpha$-consistency is $\Omega(\frac{1}{\alpha}n^{1/(\alpha-1)}\log{n})$ and our auction combines $\alpha$-consistency with $O(n^{1/(\alpha-1)}\log{n})$ robustness. Note that for any constant $\alpha$ consistency, the robustness of our auction is asymptotically optimal.
Finally, using our robustness lower bound for constant consistency, we also prove a lower bound on the “cost of smoothness” for clock auctions, i.e., a lower bound on the best achievable robustness if we also require that the performance of the auction degrades smoothly as a function of the prediction error. This result exhibits an interesting connection between consistency and smoothness and is in contrast with almost all prior work in the learning-augmented framework, which does not provide impossibility results regarding smoothness guarantees.
@@ 1.2 Related Work
Our paper lies at the intersection of two lines of literature: the literature on (deferred acceptance) clock auctions and the literature on learning-augmented algorithm and mechanism design. Below, we provide an overview of the most relevant papers in each of these two areas.
Clock auctions.
@@
Since their formal introduction in [MS2014, MS2019], clock auctions have attracted attention from both the economics and computer science communities, partly due to their many practical advantages over the sealed-bid format. These important practical properties include: (i) obvious strategyproofness, a strong notion of incentive compatibility originally defined by Li [li2017obviously]; (ii) unconditional winner privacy, defined by Milgrom and Segal [MS2019], which guarantees that clock auctions reveal the minimum possible information about the winners’ values; and (iii) transparency, including the fact that ascending clock auctions are “credible” auctions, as defined in [akbarpour2020credible]. Notably, clock auctions are the unique class of auctions satisfying these properties, making them particularly well-suited to practical implementation. We refer the reader to [MS2019] and [FGGS22, Section 1.2] for a more complete discussion of the properties of clock auctions.
@@
The papers technically closest to our work include [BLP2009, CGS22, DGR17]. Before [MS2014, MS2019] explicitly defined clock auctions, Babaioff et al. [BLP2009] proposed a deterministic, prior-free clock auction that achieved an $O(\log{v_{\max}})$-approximation, where $v_{\max}$ is the highest value of any single bidder. However, their auction notably achieved only a $\Theta(n)$-approximation. In terms of lower bounds, Dütting et al. [DGR17] demonstrated that no deterministic, prior-free clock auction could achieve a $\log^{\tau}{n}$-approximation for any constant $\tau<1$. Over a decade after [BLP2009], Christodoulou et al. [CGS22] showed that the bound in [DGR17] was essentially tight by providing a deterministic, prior-free clock auction that achieved an $O(\log{n})$-approximation (and simultaneously an $O(\log{v_{\max}})$-approximation).
In single-parameter, forward auction settings, Christodoulou et al. [CGS22] were the first to propose using randomization to circumvent the impossibility results of [DGR17]. They presented an auction that guarantees expected welfare, achieving an $O(\sqrt{\log{k}})$-approximation to the optimal social welfare, where $k$ is the number of maximal feasible sets in the set system $\mathcal{F}$. This was later improved to an $O(\log{\log{k}})$-approximation guarantee by [FGGS22].
@@
Clock auctions have also been studied beyond the single-parameter, binary service level, forward auction context. For example, [GMR17] demonstrated how to extend clock auctions to settings with multiple levels of service and analyzed several special cases of this general framework. Many works have also examined reverse (i.e., procurement) clock auctions, both with [JM17, BGGST22, ensthaler2014dynamic, BB2023, HHCT23] and without [K2015, bichler2020strategyproof] an auctioneer budget constraint. The works of [LM2020, DTR2017] studied double clock auctions involving both strategic sellers and buyers. The work of [GPPS21] studied clock auctions in the economically significant setting of interdependent values. Finally, the work of [GH23] examined clock auctions aimed at maximizing consumer utility.
Learning-augmented framework.
@@
In recent years, the learning-augmented framework has gained widespread recognition as a valuable paradigm for designing and analyzing algorithms. We refer to [MV22] for a survey of early contributions and [alps] for a frequently updated list of papers in this field. This approach aims to overcome the limitations of overly pessimistic worst-case bounds. In the past five years alone, over 200 papers have revisited classic algorithmic problems using this framework, with prominent examples including online paging [lykouris2018competitive], scheduling [PSK18], optimization problems involving covering [BMS20] and knapsack constraints [IKQP21], as well as Nash social welfare maximization [banerjee2020online], the secretary problem [AGKK23, DLLV21, KY23], and a variety of graph-related challenges [azar2022online]. As we discuss in our results section, there is another, stronger notion of consistency that requires a good approximation of the predicted solution irrespective of the quality of predictions. This notion was first proposed in [MNS07], and subsequent work, including [BMS20, KN23], also defines it as consistency. Additionally, we note that while some works do not explicitly state it, their results actually hold for this stronger notion of consistency [WZ20, JLLTZ22, ABGOT22].
@@
More closely related to our work, the line of research on learning-augmented mechanisms interacting with strategic agents is recent and was initiated by Agrawal et al. [ABGOT22] and Xu and Lu [XL22]. This line of work encompasses strategic facility location [ABGOT22, XL22, IB22, BGT24, chen2024strategic, balkanski2024randomized], strategic scheduling [XL22, BGT223], auctions [MV17, XL22, LuWanZhang23, caragiannis2024randomized, BGTZ23], bicriteria mechanism design (optimizing both social welfare and revenue) [BPS23], graph problems with private input [CKST24], metric distortion [BFGT23], and equilibrium analysis [GKST22, IBB24]. Recently, [CSV24] revisited mechanism design problems with predictions on the outcome space instead of the input. The prediction discussed in our work is one such example. For further discussion regarding this line of work, we refer the reader to [BGT23].